# STAR: A Simple Training-free Approach for Recommendations using Large Language Models

### v1

@@ Appendix A Appendix
A.1. Item Encoding Prompt for Retrieval Example
Below is an example of an item prompt for encoding with an LLM embedding API. {userbox}
@@ A.1. Item Encoding Prompt for Retrieval Example
    2. Please be aware that colors might look slightly different in person due to camera quality and monitor settings.
       Stock photos are taken in natural light with no flash.
@@
       Indicate the problem before returning. A 30-day return/exchange policy is provided as a satisfaction guarantee.
@@
A.2. Ranking Prompt Example
@@
    "title": "SHANY Professional 13-Piece Cosmetic Brush Set with Pouch, Set of 12 Brushes and 1 Pouch, Red",
@@
User: Analyze the user’s purchase history to identify user preferences and purchase patterns. Then, rank the 4 items above based on their alignment with the user’s preferences and other contextual factors. All the items should be included and listed using identifiers, in descending order of the user’s preference. The most preferred recommendation item should be listed first. The output format should be [] ¿ [], where each [] is an identifier, e.g., [1] ¿ [2]. Only respond with the ranking results, do not say any word or explain. Output in the following JSON format:

### v2

@@ 7 Limitations
Acknowledgments
We would like to thank Yupeng Hou (UCSD), Jinhyuk Lee (Google DeepMind), Jay Pujara (USC ISI), Jiaxi Tang (Google DeepMind) and Wang-Cheng Kang (Google DeepMind) for their valuable discussions and contributions to this work.
@@ Appendix A Appendix
A.1 Retrieval Performance Analysis
LLM embedding performance comparison on retrieval.
We test different LLM embedding APIs with different embedding size to understand the impact of LLM embeddings capturing item semantic similarity . Figure 6 illustrates the performance differences between the Gecko text-embedding-004 Lee et al. (2024), OpenAI text-embedding-3-small, and text-embedding-3-large models [^3] https://platform.openai.com/docs/guides/embeddings/. Overall, higher-dimensional embeddings and larger models performed better, indicating that enhanced semantic representation capabilities lead to improved semantic relationship capture.
Figure 6: Retrieval performance (Hits@50) comparison by embedding APIs for Toys and Games. The models text-embedding-004, text-embedding-3-small, and text-embedding-3-large each have a maximum dimension of 768, 1536, and 3072, respectively.
A.2 Ranking Performance Analysis
LLM performance comparison on ranking.
To evaluate how model capabilities impact ranking performance, we compare four models: gemini-1.5-flash, gemini-1.5-pro, gpt-4o-mini, and gpt-4o. In Table 8, we observe the larger models (gemini-1.5-pro and gpt-4o) tend to outperform their smaller counterparts, although the performance differences are minimal. This finding aligns with results from other studies (Qin et al., 2024), which suggest that increased model capability has limited influence on pairwise ranking tasks. Despite being more computationally expensive, pairwise ranking methods tend to be more robust than alternative approaches.
Model Beauty Toys & Games Sports & Outdoors
H@10 N@10 H@10 N@10 H@10 N@10
gemini-1.5-flash 0.101 0.060 0.120 0.073 0.056 0.034
gemini-1.5-pro 0.100 0.060 0.120 0.075 0.056 0.034
gpt-4o-mini 0.100 0.058 0.120 0.072 0.056 0.033
gpt-4o 0.100 0.060 0.121 0.074 0.056 0.033
Table 8: Pair-wise ranking performance (Hits@10 & NDCG@10) comparison by different models.
A.3 Item Encoding Prompt for Retrieval Example
Below is an example of an item prompt for encoding with an LLM embedding API. This example includes fields for title, description, category, brand, sales ranking, and price. We omit metadata fields like Item ID and URL, as those fields contain strings that can contain spurious lexical similarity (e.g., IDs: “000012”, “000013’ or URLs: “https://abc.com/uxrl”, “https://abc.com/uxrb”) and can reduce the uniformity of the embedding space and make it difficult to distinguish between semantically different items
{userbox}
@@ A.3 Item Encoding Prompt for Retrieval Example
    2. Please be aware that colors might look slightly different in person due to
       camera quality and monitor settings. Stock photos are taken in natural light with no flash.
@@ 2. Please be aware that colors might look slightly different in person due to
       Indicate the problem before returning.
       A 30-day return/exchange policy is provided as a satisfaction guarantee.
@@
A.4 Ranking Prompt Example
@@
    "title": "SHANY Professional 13-Piece Cosmetic Brush Set with Pouch,
    Set of 12 Brushes and 1 Pouch, Red",
@@
User: Analyze the user’s purchase history to identify user preferences and purchase patterns. Then, rank the 4 items above based on their alignment with the user’s preferences and other contextual factors. All the items should be included and listed using identifiers, in descending order of the user’s preference. The most preferred recommendation item should be listed first. The output format should be [] > [], where each [] is an identifier, e.g., [1] > [2]. Only respond with the ranking results, do not say any word or explain. Output in the following JSON format: